OpenStudy (ccchristian1):

Devon and his friends bought strawberry wafers for $3 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $30 to buy a total of 22 packets of wafers of the two varieties.

OpenStudy (ccchristian1):

Part A: Write a system of equations that can be solved to find the number of packets of strawberry wafers and the number of packets of chocolate wafers that Devon and his friends bought at the carnival. Define the variables used in the equations. (5 points) Part B: How many packets of chocolate wafers and strawberry wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer. (5 points)

OpenStudy (ccchristian1):

@UsukiDoll

OpenStudy (anonymous):

hello

OpenStudy (ccchristian1):

yeah?

OpenStudy (anonymous):

lets make strawberry x and chocolate y

OpenStudy (usukidoll):

oh gawd I hate these word problems sorry ^^'

OpenStudy (ccchristian1):

@UsukiDoll you got time ? lol

OpenStudy (ccchristian1):

oh alright its okay lol @UsukiDoll

OpenStudy (anonymous):

strawberry is 3 dollars so 3x chocolate is 1 dollar so 1y

OpenStudy (anonymous):

3x+y=30

OpenStudy (anonymous):

That's the first equation

OpenStudy (ccchristian1):

thats for part a?^

OpenStudy (anonymous):

the second equation is x+y=22

OpenStudy (anonymous):

both equations are part of the 'system of equations'

OpenStudy (anonymous):

For part B we need to solve for x and y.

OpenStudy (anonymous):

are you familiar with substitution?

OpenStudy (ccchristian1):

not really

OpenStudy (anonymous):

ok, so we have our two equations 3x+y=30 and x+y=22

OpenStudy (anonymous):

isolate x or y in one of the equations and then plug it into the other equation

OpenStudy (anonymous):

so isolate x in the second equation x+y=22

OpenStudy (ccchristian1):

3x+y=22?

OpenStudy (anonymous):

nope, we're just looking at x+y=22 right now

OpenStudy (anonymous):

so get x by itself

OpenStudy (ccchristian1):

so what will it be?

OpenStudy (anonymous):

we have to subtract y from both sides

OpenStudy (anonymous):

this gives x=22-y

OpenStudy (anonymous):

now we take 22-y and replace that with x in the other equation

OpenStudy (anonymous):

3(22-y)+y=30

OpenStudy (anonymous):

solve for y

OpenStudy (ccchristian1):

this will tell us how much packets did they buy

OpenStudy (anonymous):

ya

OpenStudy (ccchristian1):

then whats next ?

OpenStudy (ccchristian1):

after that

OpenStudy (anonymous):

y will tell us how many chocolate was bought. what did you get for y

OpenStudy (ccchristian1):

33?

OpenStudy (anonymous):

nope, you have to solve the equation

OpenStudy (anonymous):

first foil out 3(22-y)

OpenStudy (ccchristian1):

19?

OpenStudy (anonymous):

where are you getting these numbers?

OpenStudy (ccchristian1):

18?

OpenStudy (anonymous):

that's right, good job

OpenStudy (ccchristian1):

yeeeeeeee les get it

OpenStudy (anonymous):

so we know there are 18 chocolate wafers. how many strawberry are there?

OpenStudy (anonymous):

substitute it back into this equation x+y=22

OpenStudy (ccchristian1):

to what number ??

OpenStudy (anonymous):

well we just found that y was 18

OpenStudy (ccchristian1):

soooo......

OpenStudy (ccchristian1):

so 22?

OpenStudy (anonymous):

x+y=22 If y is 18, what is x

OpenStudy (ccchristian1):

22?

OpenStudy (anonymous):

x+(18)=22

OpenStudy (anonymous):

solve for x

OpenStudy (ccchristian1):

4?

OpenStudy (anonymous):

correct so there are 4 strawberry and 18 chocolate

OpenStudy (anonymous):

4 x 3 + 18 x 1 = 30 18 + 4 = 22

OpenStudy (anonymous):

This method is called the substitution method where you substitute one equation in for another